### v4

# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

Abstract.
High quality SQL corpus is essential for intelligent database. For example, Text-to-SQL requires SQL queries and corresponding natural language questions as training samples. However, collecting such query corpus remains challenging in practice due to the high cost of manual annotation, which highlights the importance of automatic SQL generation. Despite recent advances, existing generation methods still face limitations in achieving both diversity and cost-effectiveness. Besides, many methods also treat all tables equally, which overlooks schema complexity and leads to under-utilization of structurally rich tables. To address these issues, this paper proposes a multi-agent framework for high-quality and large-scale SQL generation, dubbed SQL-Factory. It decomposes the generation process into three collaborative teams: the Generation Team explores diverse query structures using a powerful language model, the Expansion Team scales promising patterns via a lightweight language model, and the Management Team adaptively schedules the workflow and evaluates the quality of synthesized queries. This modular framework ensures a balanced trade-off between diversity, scalability, and generation cost. We apply SQL-Factory to four widely used benchmarks and generate over 300,000 SQL queries with less than \$200 API cost. Our generated queries achieve higher diversity compared to other methods, and extensive experiments demonstrate that the generated queries significantly improve the model performance in various downstream tasks.
Yuren Mao is the corresponding author.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/LJHzju/SQL-Factory.
1. Introduction
Figure 1. A conceptual illustration of SQL-Factory’s multi-agent framework. The framework consists of six agents: (1) Table Selection Agent, (2) Generation Agent, (3) Management Agent, (4) Critical Agent, (5) Seed Selection Agent and (6) Expansion Agent.
A high-quality SQL corpus is essential for advancing intelligent database technologies, such as cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023), configuration tuning (Zhang et al., 2019; Li et al., 2025b; Van Aken et al., 2017; Bianchi et al., 2024) and natural language interfaces (Giaquinto et al., 2023; Li et al., 2024; Sun et al., 2023). For example, to train a learning-based configuration tuning system, it is necessary to observe how different SQL queries interact with various configuration settings, which requires a large number of diverse SQL workloads. Besides, since SQL can be easily translated into natural language, it also facilitates the construction of datasets for tasks like Text-to-SQL.
However, obtaining such corpus in practice remains highly challenging. The process typically requires labor-intensive manual annotation by domain experts, resulting in high costs in terms of time and resources. Meanwhile, publicly available SQL corpora are scarce, as real-world workloads are often restricted due to commercial sensitivity, or limited access to production environments. Therefore, automatic SQL generation has gained increasing attention (Zhang et al., 2022; Ding et al., 2021). Prior research has explored a variety of techniques, including (1) Template-based methods, which rely on hand-crafted query templates to produce interpretable yet structurally limited SQL instances (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018); (2) Grammar-guided random generation, which constructs queries by sampling from SQL grammar rules but often lacks semantic coherence (Seltenreich, 2016; Slutz, 1998; Abdul Khalek and Khurshid, 2010); and (3) Learning-based approaches, which employ machine learning models to generate queries in a data-driven manner (Zhang et al., 2022; Liu et al., 2018; Li et al., 2025a). However, existing methods often struggle to simultaneously achieve sufficient query diversity, schema-aware control, and cost-effective generation.
Insufficient query diversity. A lack of diversity means that the generated SQL queries tend to follow similar structures, use the same types of operators, or involve limited combinations of joined tables. Template-based approaches such as TPC-DS (Nambiar and Poess, 2006) and Join Order Benchmark (JOB) (Leis et al., 2018) typically rely on fixed syntactic forms, which restrict their ability to produce varied query types. Learning-based methods such as TreeGAN (Liu et al., 2018) and LearnedSQLGen (Zhang et al., 2022) exhibit strong tendencies to overfit to frequent query skeletons in training corpora, producing workloads clustered around recurrent table subsets and operator chains. As a result, these methods fail to capture the rich variety of queries that appear in real-world workloads. Improving query diversity would help cover more edge cases and make trained models more robust in practical database environments.
Lack of schema-aware planning. Effective SQL generation should account for schema heterogeneity by selectively targeting complex or underutilized tables. For instance, tables with many columns or dense foreign key connections are typically involved in more intricate query logic and should therefore receive more attention during workload construction. However, most existing methods treat all tables equally, applying uniform sampling that ignore these schema-level differences. Random generation methods such as SQLSmith (Seltenreich, 2016) sample syntax trees without considering schema semantics or historical usage patterns, leading to generating queries that either overfit frequent tables or neglect structurally rich regions of the schema. Although OmniSQL (Li et al., 2025a) incorporates schema context into its prompts, it does not explicitly guide table selection based on schema-level statistics. This lack of adaptive planning limits its ability to ensure balanced coverage or prioritize structurally informative parts of the schema.
High generation cost. Existing methods based on large language models (LLMs) often depend on either proprietary APIs or heavyweight open-source models to achieve reliable quality, leading to high costs in API usage or deployment. For instance, OmniSQL (Li et al., 2025a) demonstrates the feasibility of large-scale SQL generation by using a mixture of locally deployed models, including those with over 30 billion parameters. This approach significantly increases deployment costs.
To address these limitations, we introduce SQL-Factory, a multi-agent framework designed for high-quality and large-scale SQL query generation. SQL-Factory leverages three collaborative agent teams: the Management Team, the Generation Team, and the Expansion Team, each responsible for a distinct aspect of the generation process. The Generation Team explores structurally diverse and expressive SQL patterns using a powerful large-scale language model, ensuring sufficient variety across query forms and structures. The Expansion Team complements this by scaling the workload through a lightweight and cost-effective local model, which generate semantically coherent variants based on representative seed queries, significantly reducing reliance on expensive models. Finally, the Management Team evaluates query quality and schema-level distribution in real time, guiding scheduling decisions to ensure that structurally complex or underrepresented schema components are adequately covered. This division of responsibilities enables SQL-Factory to achieve a promising trade-off between query diversity, generation efficiency, and schema-level balance.
In summary, the key contributions of this paper can be outlined as follows:
  • We propose SQL-Factory, a multi-agent SQL generation framework that introduces three specialized teams for generation, expansion, and management, enabling high-quality and large-scale SQL synthesis.
  • We design a self-correction generation workflow that dynamically integrates structural exploration (Generation) and pattern exploitation (Expansion). The Management Team adaptively schedules different agents based on real-time evaluation of query quality.
  • We generate over 300,000 broadly distributed SQL queries across four widely used benchmarks: TPC-DS, IMDB, BIRD, and SPIDER. Extensive experiments demonstrate that these queries exhibit significantly higher structural and semantic diversity compared to existing generation methods, and further improve performance on downstream tasks such as Text-to-SQL modeling and SQL clustering.
2. Related Work
2.1. Data-Driven Intelligent Database
The integration of data-driven artificial intelligence techniques, which leverage models trained on large-scale datasets, into database systems has emerged as a rapidly advancing research area. This paradigm explores the use of learning-based approaches, particularly large language models, to improve various database-related tasks such as configuration optimization, cardinality estimation, and Text-to-SQL translation.
One major area of research focuses on database configuration tuning. Various automatic tuning methods (Van Aken et al., 2017; Zhang et al., 2019; Li et al., 2025b) have been proposed to recommend high-performance system configurations, thereby reducing manual effort and improving overall throughput. These methods typically rely on learning from historical workload logs to inform future configuration decisions.
Another important line of work focuses on converting natural language to SQL (NL2SQL). A wide range of methods (Pourreza and Rafiei, 2023; Wang et al., 2025b; Gao et al., 2024; Zhang et al., 2024; Xu et al., 2025; Dong et al., 2023) train neural models on aligned natural language and SQL query pairs, enabling users to express complex information needs in plain language and receive executable queries as output.
Beyond these areas, research in this domain has also extended to cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023), index selection (Yu et al., 2024; Wang et al., 2025a), and query clustering. These applications benefit from the pattern recognition and generalization abilities of modern neural models.
Despite these advancements, most of them rely on the availability of high-quality SQL corpus for training or simulation. However, obtaining such a corpus is highly challenging in practice due to the labor-intensive and costly manual annotation by domain experts familiar with SQL and schema semantics, as well as the limited availability of real-world queries. This creates a growing need for automatic SQL generation frameworks.
2.2. SQL Generation
SQL generation aims to automatically synthesize large volumes of semantically meaningful and structurally diverse queries tailored to a specific database schema or system. This task plays an essential role in supporting downstream applications such as system benchmarking, robustness testing, and data-driven model training.
Early methods for SQL generation were primarily rule-based or template-driven. For example, industry-standard benchmarks such as TPC-DS (Nambiar and Poess, 2006), TPC-H (Poess and Floyd, 2000) and Join Order Benchmark (JOB) (Leis et al., 2018) construct queries from pre-defined templates, offering a controllable and interpretable means of workload construction. However, due to the limited number of templates, these methods often struggle to cover diverse query patterns and exhibit limited query randomness or structural variability. Random generation methods (Seltenreich, 2016; Slutz, 1998; Abdul Khalek and Khurshid, 2010) attempt to increase diversity by randomly walking through SQL grammar rules or syntax trees. These methods increase lexical diversity but lack semantic grounding, often producing unexecutable or meaningless queries that fail to reflect realistic workload patterns.
Recent research has explored the use of machine learning and large language models for SQL generation. For example, TreeGAN (Liu et al., 2018) proposed a GAN-based framework that synthesizes SQL queries by imitating the structure of existing workloads. However, it requires large volumes of representative SQL data for training and struggles to generalize beyond seen patterns or schemas. LearnedSQLGen (Zhang et al., 2022) introduces a reinforcement learning, aiming to synthesize queries that satisfy specific execution constraints learning approach that generates queries under execution constraints such as cardinality or cost. However, it often lacks structural diversity, as the model tends to exploit familiar query patterns that satisfy the constraints rather than exploring novel or complex structures. OmniSQL (Li et al., 2025a) adopts a large-scale data generation pipeline that synthesizes millions of queries using a set of locally employed models. While it achieves impressive scale, the framework relies heavily on high-cost models and lacks intermediate control during query synthesis. Without mechanisms to guide schema usage or enforce diversity, OmniSQL often produces redundant queries and cannot adapt generation focus in real time. These limitations reduce its efficiency, especially under resource constraints.
2.3. LLM-based Multi-Agent System
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and reasoning, enabling them to perform a wide range of complex tasks with human-like proficiency. To organize these capabilities effectively, single-agent architectures typically equip an LLM with modular components such as Profile, Memory, Planning, and Action (Xi et al., 2025), which allow it to decompose tasks, retrieve relevant information, and simulate multi-step workflows. This design has proven effective in various domains, including dialogue generation, retrieval-augmented generation, and structured data interaction (Wang et al., 2024; Modarressi et al., 2023; Hu et al., 2023).
Building upon this foundation, recent research has extended LLM capabilities into collaborative multi-agent systems, where multiple specialized agents coordinate to complete tasks through dynamic communication and role-based reasoning. Multi-agent systems distribute cognitive responsibilities across distinct agents, each tailored for specific subgoals or expertise areas. This paradigm fosters more interpretable, scalable, and resilient behavior, particularly in settings that demand iterative planning, task delegation, and adaptive feedback handling. Multi-agent LLM systems have shown promise in domains such as software engineering (Qian et al., 2024; Chen et al., 2023), where agents representing roles like developer, tester, and reviewer to build and refine software artifacts. Similarly, in scientific experimentation (Zheng et al., 2023) and medical diagnostics (Tang et al., 2024), these systems enable complex reasoning chains by leveraging role specialization and interactive deliberation among agents.
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more robust and accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage several large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
3. System Overview
Figure 2. An overview of SQL-Factory.
This section introduces SQL-Factory, a schema-aware and cost-efficient multi-agent framework for automatically synthesizing large-scale and diverse SQL query workloads. As shown in Figure 2, SQL-Factory consists of three collaborative agent teams: (1) the Management Team, (2) the Generation Team, and (3) the Expansion Team. These teams work together to address three key challenges in SQL generation, including insufficient structural diversity, lack of schema-aware planning, and high generation cost. In addition, SQL-Factory maintains a centralized SQL Pool that stores all synthesized queries throughout the process.
Generation Team. The Generation Team is designed to maximize structural diversity of generated queries, tackling the problem of repetitive or overly templated query outputs. It comprises the Table Selection Agent and the Generation Agent. The Table Selection Agent realizes schema-aware planning by prioritizing structurally complex tables or those with few synthesized queries. Then, the Generation Agent employs a powerful LLM to generate novel and semantically valid queries for selected tables. These queries often include complex structures such as nested subqueries, aggregations, and multi-table joins. By focusing on pattern exploration, the Generation Team ensures the introduction of high-diversity query structures that guide subsequent expansion.
Expansion Team. The Expansion Team addresses the cost-efficiency challenge by scaling up the number of queries using lightweight generation. It includes the Seed Selection Agent and the Expansion Agent. The Seed Selection Agent identifies high-quality queries that involve under-covered tables, and the Expansion Agent uses them as seed queries to synthesize new variants via controlled augmentation. The Expansion Team concentrates on scaling promising patterns to achieve high query volume with minimal computational overhead, enabling the system to achieve both depth and breadth in SQL generation.
Management Team. The Management Team serves as the central controller of SQL-Factory, responsible for coordinating these two teams and improving overall quality by filtering out invalid or overly similar queries. It plays a crucial role in maintaining system efficiency and adaptability, directly addressing the challenge of balancing structural diversity with generation cost. This team consists of two components: the Critical Agent and the Management Agent. The Critical Agent performs real-time quality assessments using an Execution Tool to check query validity and a Retrieval Tool to estimate semantic redundancy via hybrid similarity. Based on these evaluations, the Management Agent dynamically switches between exploration (Generation Team) and exploitation (Expansion Team) to balance novelty and efficiency.
SQL Pool. The SQL Pool is the centralized repository for all generated and expanded SQL queries. Implemented with a relational database such as MySQL, it is composed of two tables: Table $R$, which stores the detailed SQL query along with the list of referenced tables, and Table $E$, which maintains precomputed embedding vectors for similarity measurement and retrieval tasks.
4. methodology
In this section, we provide a detailed methodology for SQL-Factory by describing the internal design and its key components, including the Management Team, Generation Team and Expansion Teams.
4.1. Management Team
The Management Team is responsible for orchestrating the entire SQL synthesis pipeline by coordinating interactions between the Generation Team and the Expansion Team. Its core objective is to maintain a balanced trade-off between structural diversity and data scalability while ensuring that only high-quality SQL queries are retained. This team consists of two components: the Management Agent, which governs the overall coordination of SQL synthesis, and the Critical Agent, which evaluates the quality of SQL queries from multiple perspectives.
4.1.1. Management Agent
The Management Agent governs the system state by dynamically selecting whether the Generation Team or Expansion Team should be invoked in each round. Its goal is to adaptively balance between generating structurally novel queries (exploration) and expanding promising query patterns based on existing queries (exploitation). This agent is implemented using a large language model, enabling flexible reasoning over both system statistics and quality metrics obtained from the Critical Agent.
Formally, the scheduling process is modeled as a discrete state transition function:
(1) $S_{t+1}=\text{Agent}_{\text{M}}(S_{t},E_{t}),$
where $S_{t}\in[\text{GEN},\text{EXP}]$ denotes the current system state (i.e., invoking the Generation or Expansion Team), and $E_{t}$ represents the quality evaluation result provided by the Critical Agent at round $t$, including metrics such as query executability and average structural similarity. The policy function $\text{Agent}_{\text{M}}(\cdot)$ is implemented via LLM-based inference, which selects the next state $S_{t+1}$ based on statistical trends observed in the evaluation history.
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for structurally complex tables or those with few synthesized queries. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from the Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that remains both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns.
4.1.2. Critical Agent
The Critical Agent serves as the evaluation engine of SQL-Factory. Its primary function is to assess the quality of newly generated or expanded SQL queries on the fly before they are stored into the SQL Pool. This agent operates as a shared component used by both the Generation Team and the Expansion Team, and its evaluation result also plays a crucial role in guiding the Management Agent’s scheduling decisions.
Specifically, the Critical Agent comprises two modular components: the Execution Tool $\text{Tool}_{\text{E}}$, which verifies query executability, and the Retrieval Tool $\text{Tool}_{\text{R}}$, which using hybrid similarity to retrieve the most similar SQL queries. The evaluation process differs slightly depending on whether the query was produced by the Generation Team or the Expansion Team.
For SQL queries from the Generation Team, the powerful large model ensures a degree of novelty and exploratory coverage. Therefore, we focus primarily on syntactic and semantic validity. Given a set of generated queries $\mathcal{Q}_{t}=\{q_{t}^{1},q_{t}^{2},\dots,q_{t}^{n}\}$ at step $t$, the Critical Agent call the Execution Tool $\text{Tool}_{\text{E}}$ to compile each query on the target database schema and reports a binary executability flag. The overall evaluation for this batch is defined as:
(2) $E_{t}=\text{Agent}_{\text{C}}(\mathcal{Q}_{t},\text{Tool}_{\text{E}}),$
Figure 3. An overview of our hybrid similarity.
In contrast, queries from the Expansion Team are subject to stricter evaluation to mitigate semantic redundancy. Each query is first passed through the Execution Tool as above, and then evaluated with the help of the Retrieval Tool. The Retrieval Tool encodes each query into a dense embedding vector using a pretrained language model. It then performs approximate nearest neighbor (ANN) retrieval to collect the top-$(k\times 20)$ candidates from the SQL Pool based on cosine similarity. From the retrieved candidate set, we select the top-$k$ most relevant queries for final evaluation using a hybrid similarity metric.
We introduce this hybrid metric to address the inherent limitations of embedding-based cosine similarity, which primarily captures surface-level semantic proximity. While effective for general clustering tasks, cosine similarity often fails to capture deeper structural or logical differences between SQL queries that have similar tokens but differ significantly in their execution semantics or syntax trees. To produce a more robust and semantically grounded similarity measure, we integrate multiple perspectives into a unified scoring framework. An illustration of the hybrid similarity computation is provided in Figure 3.
Formally, the similarity between a candidate query $q$ and a historical query $q^{\prime}\in\text{SQL Pool}$ is computed as:
(3) $\text{Sim}(q,q^{\prime})=\alpha\cdot\text{Sim}_{\text{Tok}}+\beta\cdot\text{Sim}_{\text{AST}}+\gamma\cdot\text{Sim}_{\text{Emb}},$
where $\alpha$, $\beta$, and $\gamma$ are the weights assigned to the three components. Empirically, we set $\alpha=0.6$, $\beta=0.3$, and $\gamma=0.1$ to balance between surface-level lexical overlap, deeper structural and semantic alignment. Each component here captures a different aspect of query similarity:
  • Token-level similarity $\text{Sim}_{\text{Tok}}$ is computed via normalized token to measure lexical alignment after token sorting.
  • AST-level similarity $\text{Sim}_{\text{AST}}$ evaluates structural similarity based on the abstract syntax trees (ASTs) of the two SQL queries, defined as:
    (4) $\text{Sim}_{\text{AST}}(q,q^{\prime})=1-\frac{\text{EditDist}(\text{AST}(q),\text{AST}(q^{\prime}))}{\max(|\text{AST}(q)|,|\text{AST}(q^{\prime})|)},$
    where $\text{EditDist}(\cdot)$ expresses the edit distance between AST trees, and ||$\cdot$|| ⋅ | represents the total number of nodes in the AST. This reflects the structural overlap between query patterns, including shared clauses and logical operators.
  • Embedding-level similarity $\text{Sim}_{\text{Emb}}$ is computed using cosine similarity between dense representations. It captures semantic alignment in latent space, abstracting away surface-level variations, defined as:
    (5) $\text{Sim}_{\text{Emb}}(q,q^{\prime})=\frac{\text{Emb}(q)\cdot\text{Emb}(q^{\prime})}{||\text{Emb}(q)||\ ||\text{Emb}(q^{\prime})||},$
    where $\text{Emb}(q)$ denotes the vector representation of query $q$ encoded by the BERT encoder, and |||||$\cdot$||| | ⋅ | | represents the norm of the vector.
By integrating the above three similarity metrics, this hybrid scoring function provides a more comprehensive assessment of query similarity. It is particularly effective in identifying subtle redundancies among expanded SQL queries.
Finally, the Critical Agent aggregates the execution validation results from $\text{Tool}_{\text{E}}$ and retrieval results from $\text{Tool}_{\text{R}}$ to produce the final evaluation score.
(6) $E_{t}=\text{Agent}_{\text{C}}(\mathcal{Q}_{t},\ \text{Tool}_{\text{E}},\ \text{Tool}_{\text{R}}),$
The final evaluation result $E_{t}$ is then passed to the Management Agent to support real-time scheduling decisions.
4.2. Generation Team
The Generation Team is responsible for initiating new SQL patterns through schema-aware query construction. This team is activated when the system enters the exploration phase, typically after the Management Agent detects either insufficient diversity or excessive redundancy in the current SQL Pool. The primary goal of the Generation Team is to expand the search space by synthesizing novel and complex queries, thereby enriching the structural and logical variety of the dataset.
Figure 4. Generation Prompt Template.
To achieve this, the Generation Team operates in a two-step pipeline. First, the Table Selection Agent analyzes schema-level statistics from the SQL Pool to identify underrepresented tables that have not been sufficiently queried. Then, the Generation Agent uses a large language model to construct diverse SQL queries over the selected tables. These queries are generated from scratch, without direct reference to existing SQL patterns, ensuring minimal bias and maximal novelty.
By strategically prioritizing structurally rich but underutilized schema components and leveraging the expressive power of LLMs, the Generation Team enables SQL-Factory to systematically explore the vast query space and discover new SQL structures that are unlikely to emerge from pattern-based expansion alone.
4.2.1. Table Selection Agent
The Table Selection Agent is responsible for identifying which tables should be prioritized for SQL generation during the exploration phase. Its goal is to maximize structural diversity by focusing on underrepresented or structurally rich tables in the database. This agent forms the first step in the generation pipeline and directly influences the quality and novelty of queries synthesized by the subsequent Generation Agent.
The agent operates over two sources of information: (1) the full schema of the target database, including all table names and their respective column types; and (2) a statistical summary of SQL usage derived from the SQL Pool, such as the number of queries for each table. These metrics are extracted and structured via a dedicated Statistical Tool $\text{Tool}_{S}$, which computes table-level query coverage.
Formally, let $\mathcal{T}=\{T_{1},T_{2},\dots,T_{m}\}$ denote the complete set of tables in the current database. The full input to the Table Selection Agent is then:
(7) $\mathcal{S}=\left\{<\text{Schema}(T_{i}),\;\text{Tool}_{\text{S}}(\text{SQL Pool},\ T_{i})>\;\middle|\;T_{i}\in\mathcal{T}\right\},$
The LLM is then prompted with $\mathcal{S}$ to select a subset of tables $\mathcal{T}^{\prime}\subset\mathcal{T}$ that are sparsely covered. Formally, the output of the Table Selection Agent is defined as:
(8) $\mathcal{T}^{\prime}=\text{Agent}_{\text{T}}(\mathcal{S}),\quad\text{where}\ \mathcal{T}^{\prime}\subset\mathcal{T}$
By integrating schema-level complexity and query statistics, the Table Selection Agent enables informed decision-making that balances exploration of new structural patterns with targeted improvement of under-sampled schema regions. This modular design also allows the system to dynamically adapt its generation focus based on the evolving distribution of queries in the SQL Pool.
4.2.2. Generation Agent
The Generation Agent is responsible for synthesizing new SQL queries in a schema-aware and open-ended manner. It takes as input the set of tables $\mathcal{T}^{\prime}$ selected by the Table Selection Agent and aims to produce queries that are structurally diverse. This process constitutes the core exploration step in SQL-Factory, enabling the discovery of novel query patterns that cannot be easily obtained by simply modifying existing SQL queries.
The agent is implemented using a powerful, instruction-tuned large language model, which is prompted with both schema information for $\mathcal{T}^{\prime}$ and high-level generation instructions, as shown in Figure 4. Rather than treating each table independently, the model considers the joint schema of all selected tables and is encouraged to synthesize complex queries that include multi-table joins, nested subqueries, aggregations, and window functions. Besides, to guide the model effectively, the schema information included in the prompt is carefully constructed to ensure both syntactic validity and semantic meaningfulness. For each selected table, we encode not only its column names and data types, but also relevant content-aware hints to reduce the likelihood of generating vacuous or logically invalid queries. Specifically, for columns of enumerated types (e.g., status, category), we provide up to five representative values from the domain to help the model infer plausible filter conditions or join constraints. For numeric or date type columns (e.g., int, float, date), we include its minimum and maximum values in the database, offering statistical context for range-based predicates (e.g., price ¿ 100). For free-text or string columns, we randomly sample up to five values as textual examples, assisting the model in generate meaningful LIKE conditions or equality conditions. These content-aware augmentations are inserted inline with the schema description in a structured format, forming an enriched prompt that provides the model with both structural knowledge and semantic information. This design significantly improves the quality and contextual relevance of the generated queries, particularly for tables with sparse usage history or rich attribute semantics. Finally, the output of the Generation Agent are summarized as:
(9) $\mathcal{Q}_{\text{gen}}=\text{Agent}_{\text{G}}(\mathcal{T}^{\prime})=\{q_{1},q_{2},\dots,q_{k}\}.$
The generation is carried out in an open-ended fashion, without referencing queries from the SQL Pool. To ensure structural richness, prompts are explicitly designed to avoid trivial query types such as single-table scans or simple projections. This setup allows the Generation Agent to push the boundaries of the query space and introduce SQL patterns that improve structural diversity.
All generated queries $\mathcal{Q}_{\text{gen}}$ are forwarded to the Critical Agent for executability validation. Only those that compile successfully are inserted into the SQL Pool. In this way, the Generation Agent supports horizontal dataset expansion, discovering new structural dimensions of the underlying database schema.
4.3. Expansion Team
Figure 5. Expansion Prompt Template.
The Expansion Team is designed to scale up the SQL dataset by exploiting existing query structures through controlled augmentation. This team is activated during the exploitation phase, particularly after the Management Agent determines that the diversity of the SQL Pool has been sufficiently enriched and the system should focus on increasing query volume. Unlike the Generation Team, which relies on a large-scale language model to generate structurally novel queries from scratch, the Expansion Team utilizes a lightweight model to synthesize new SQL queries by learning from high-quality seeds already present in the SQL Pool.
The expansion process also follows a two-step workflow. First, the Seed Selection Agent identifies representative queries from the SQL Pool that involve under-covered tables or patterns. Then, the Expansion Agent learns from these queries and generates structurally similar but lexically and semantically varied SQL queries that preserve executability. Since this phase focuses on local exploration around known patterns, it enables SQL-Factory to rapidly increase the number of high-quality queries while maintaining system efficiency. Together, these two agents complements the structural exploration performed by the Generation Team and allows SQL-Factory to produce a large volume of diverse and executable queries at scale with low cost.
4.3.1. Seed Selection Agent
The Seed Selection Agent plays a similar role to the Table Selection Agent in the Generation Team, as both are responsible for identifying underrepresented schema components for further query synthesis. However, rather than initiating query generation from scratch, the Seed Selection Agent focuses on retrieving representative queries from the existing SQL Pool to serve as seeds for the Expansion Agent. These seed queries act as structural blueprints, enabling the system to scale existing patterns while preserving semantic integrity.
To select appropriate seeds, the agent first determines a set of underutilized tables $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ by analyzing query distribution statistics via a shared statistical module $\text{Tool}_{\text{S}}$. It then search the SQL Pool to collect a small set of seed queries $\mathcal{Q}_{\text{seed}}$, each of which must reference at least one table from $\mathcal{T}^{\prime}$. These queries are selected not only based on table inclusion but also for their structural validity and potential for lexical or syntactic variation. This approach allows the Expansion Agent to operate in a pattern-aware and data-driven manner, leveraging past structures to efficiently generate new and diverse SQL variants. Formally, the the final output of Seed Selection Agent is:
(10) $\mathcal{Q}_{\text{seed}}=\text{Agent}_{\text{S}}(\mathcal{T}^{\prime})=\{q_{i}\in\mathcal{Q}_{\text{pool}}\mid\text{Tables}(q_{i})\cap\mathcal{T}^{\prime}\neq\emptyset\}.$
By identifying high-quality yet underutilized queries, the Seed Selection Agent enables the Expansion Team to efficiently enrich the dataset without redundant exploration.
4.3.2. Expansion Agent
The Expansion Agent generates new SQL queries by leveraging representative seed queries retrieved by the Seed Selection Agent. Unlike the Generation Agent, which synthesizes queries from scratch based on schema information, the Expansion Agent focuses on local exploration. Its goal is to generate syntactically diverse and executable queries by learning from the structure or intent of seed queries. This agent enables the system to scale up query volume without incurring the computational cost of large-scale model inference.
Specifically, this agent operates on the set of seed queries, which were selected based on their relevance to underrepresented tables. It uses a lightweight model to derive multiple variants based on these seed queries, employing techniques such as structural rewriting, substructure recombination. Besides, the agent is not limited to structural transformations but also introduces lexical and logical variations to ensure semantic diversity.
As shown in Figure 5, each prompt includes a small set of seed SQL queries and task-specific instructions. These instructions encourage the model to generate structurally diverse and executable variants, while inferring schema constraints implicitly from the provided queries. Formally, given a set of seed queries $\mathcal{Q}_{\text{seed}}$, the Expansion Agent generates a corresponding set of SQL variants $\mathcal{Q}_{\text{exp}}$ as follows:
(11) $\mathcal{Q}_{\text{exp}}=\text{Agent}_{\text{E}}(\mathcal{Q}_{\text{seed}}).$
After expansion, the resulting queries $\mathcal{Q}_{\text{exp}}$ are passed to the Critical Agent. In this phase, queries are not only checked for executability but also evaluated for semantic redundancy using retrieval-based similarity assessment. Only queries that meet the criteria for both executability and novelty are retained and added to the SQL Pool. Through this process, the Expansion Agent supports vertical dataset expansion. It strengthens the density of high-quality query patterns within the local neighborhood of schema space, ultimately improving the dataset’s expressiveness and utility in downstream tasks. While the Generation Agent explores new structural forms, the Expansion Agent intensifies local pattern regions, ensuring both diversity and volume.
5. Experiments
In this section, we conduct extensive experiments to demonstrate the efficiency of the SQL-Factory framework from multiple perspectives, including detailed analysis on query quality, and its effectiveness on downstream tasks.
5.1. Experiment Setup
5.1.1. Environment
We conduct the experiments on one GPU servers, equipped with 4 NVIDIA RTX-4090-24GB GPUs, an Intel(R) Xeon(R) Silver 4316 CPU, and 256GB of RAM. For LLM training, we utilize PyTorch 2.3.0, DeepSpeed 0.14.0 and LLaMA-Factory 0.8.4, while vLLM 0.4.3 is employed for LLM inference.
5.1.2. Benchmarks
To evaluate the effectiveness and generality of SQL-Factory, we select four typical database benchmarks for synthesizing queries. (1) TPC-DS (Nambiar and Poess, 2006) is an industry-standard benchmark for decision support systems, featuring a complex schema with 24 tables and various star-join patterns. It serves as a rigorous testbed for generating analytical SQL queries. (2) IMDB (Maas et al., 2011) is a movie-related database, which includes 21 tables on movies and actors. Its real-world structure and moderate size make it suitable for generating realistic join-heavy queries. (3) SPIDER (Yu et al., 2018) is a large-scale text-to-SQL dataset consisting of several databases with highly diverse and normalized schemas. We use the SPIDER schema collection as a target set for multi-domain query generation. (4) BIRD (Li et al., 2023a) is a recent benchmark introduced to evaluate database-instruction reasoning in realistic and complex environments. We use BIRD both as a target for SQL generation and as a downstream testbed for Text-to-SQL evaluation, particularly to assess robustness under schema ambiguity. Across all benchmarks, we parse the schema into structured representations including table names, column types, and foreign key relations. These are used as inputs for SQL-Factory’s Generation and Expansion workflows.
Besides, To evaluate the effectiveness of the generated SQL queries in improving representation learning, we adopt three widely-used query clustering benchmarks: IIT Bombay (Chandra et al., 2015), UB Exam, and PocketData (Kennedy et al., 2016). These datasets contain real-world SQL queries collected from academic and industry environments and their manually annotated cluster labels.
5.1.3. Baseline
We organize our baselines into three categories to cover different stages of evaluation: SQL generation, Text-to-SQL, and SQL query clustering.
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both random generation methods and learning-based generation methods. Specifically, we include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
Text-to-SQL Baselines. For evaluating the effectiveness of the generated data on downstream Text-to-SQL tasks, we fine-tune two base models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, using SQL-NL pairs derived from the SQL queries synthesized by SQL-Factory. The NL annotations are obtained via a post-processing step, as described in Section 5.3. We compare the performance of these fine-tuned models with several model of similar scale, including Codes-1B, and Codes-3B (Li et al., 2024). This setup allows us to examine whether augmenting training with generated queries generated by SQL-Factory can improve model performance.
Query Clustering Baselines. For the SQL clustering task, we evaluate several established similarity models. These include statistical and heuristic approaches such as Aouiche (Aouiche et al., 2006), Aligon (Aligon et al., 2014), and Makiyama (Makiyama et al., 2015), the pretrained model-based PreQR, and two general-purpose language models: BERT-Base and BERT-Large (Devlin et al., 2019). We further include versions of BERT fine-tuned with contrastive learning on SQL-Factory data to highlight the effectiveness of our generated queries in improving SQL representation learning.
5.1.4. Implementation Details
We utilize different large language models and configurations to implement these agents in SQL-Factory. Each module is configured to balance quality and efficiency.
Generation Agent We adopt GPT-4o (OpenAI, 2024) as the backbone model for the Generation Agent, due to its strong capabilities in zero-shot reasoning and diverse SQL synthesis. The model is accessed through the OpenAI API, with generation configured using a temperature of 0.8 and top-$p$ sampling for better output diversity.
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following capability. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
Other Agents. The Management Agent, Critical Agent, Table Selection Agent, and Seed Selection Agent also reused Qwen2.5-Coder-14B to reduce deployment overhead. Here we use a lower temperature of 0.3, aiming for more deterministic and stable reasoning behavior without sacrificing overall coherence.
5.2. Quality Analysis of Generated SQL
We apply SQL-Factory to synthesize large-scale SQL queries across the four benchmarks. Each benchmark varies in schema size and complexity, enabling us to test the scalability and adaptability of our multi-agent generation framework.
5.2.1. Overview of Generated Queries
((a))
((b))
((c))
((d))
((e))
Figure 6. Overview of SQL queries synthesized by SQL-Factory.
Key aspects such as join depth, predicate usage, subquery nesting, aggregate usage, and overall token length provide meaningful insights into the expressiveness and structural diversity of SQL queries. These dimensions serve as proxies for measuring the coverage of different logical constructs and the capability to reflect real-world query distributions. Therefore, we conduct a comprehensive analysis of the SQL queries along these dimensions in Figure 6.
Figure 6(a) illustrates the distribution of join operations per query. Notably, 78.6% of all queries involve at least one join, and 30% contain two or more join. This indicates that SQL-Factory effectively generates multi-table queries, showcasing its capacity for relational reasoning and schema-aware generation. Figure 6(b) analyzes the number of predicates in queries. More than 58% of the generated queries contain two or more predicates, with 9.8% having four or more. This confirms that SQL-Factory not only generates structurally rich queries but also embeds fine-grained filtering logic, which is critical for supporting meaningful data retrieval and realistic workload simulation. Besides, Figure 6(c) and Figure 6(d) present the proportions of queries that contain nested subqueries and aggregation operations, respectively. These results demonstrate that a large portion of the generated SQLs utilize advanced SQL constructs, further validating the model’s ability to produce structurally rich and expressive queries. Finally, Figure 6(e) reports the distribution of token lengths across these queries. The x-axis represents the number of tokens for each query, while the y-axis shows the corresponding frequency. The distribution reveals a broad range of query lengths, confirming that SQL-Factory does not produce templated or overly uniform queries, but rather supports diverse query formulations at varying levels of complexity.
Overall, these results confirm that SQL-Factory produces syntactically diverse and structurally expressive SQL queries, with rich combinations of logical constructs. Such diversity is essential for training robust models in downstream applications.
5.2.2. Cost Analysis of SQL Generation
Table 1. The statistics of the generated SQL.
Benchmark # DB # Table # Col/DB # Queries Price (\$)
TPC-DS (Nambiar and Poess, 2006) 1 24 449 50,000 51.78
IMDB (Maas et al., 2011) 1 21 108 40,000 18.47
Spider (Yu et al., 2018) 69 795 26.82 130,000 42.16
BIRD (Li et al., 2023a) 146 524 54.56 160,000 67.50
Table 1 summarizes the key statistics of SQL queries generated by SQL-Factory across four benchmarks. These include the number of databases, tables, average number of columns per database, the total number of generated queries, and the corresponding cost of API.
The four benchmarks span a diverse range of schema structures. TPC-DS and IMDB each consist of a single database but vary substantially in complexity, TPC-DS features 449 columns while IMDB has 108 specifically. Meanwhile, BIRD and Spider contain hundreds of databases with fewer columns per schema on average, representing more fragmented but wider coverage scenarios. Across these benchmarks, we observe that the total number of generated queries generally correlates with the schema’s complexity and scale. For example, the largest numbers of queries are generated for BIRD (160,000) and Spider (130,000), which contain the most tables and databases. In contrast, IMDB yields 40,000 SQLs due to its simpler, single-schema structure. This adaptive generation behavior emerges naturally from our multi-agent framework, which allocates more generation rounds and expansions in response to richer or more varied schema coverage requirements.
Despite the varying generation needs, SQL-Factory maintains high efficiency. The total cost across all four benchmarks remains under \$200. Notably, if the entire workload were generated solely by GPT-4o without the Expansion Team, the API cost would approach \$1000, highlighting the substantial cost savings enabled by our hybrid design. Together, these results demonstrate that SQL-Factory is capable of producing SQL workloads at scale, while maintaining both schema adaptivity and cost-effectiveness.
5.2.3. Comparative Analysis on Query Diversity
Table 2. Comparison with Existing Query Generation Methods in Terms of Diversity.
Method Hybrid Vendi Score Vendi Score
Similarity SQLEncoder SimCSE
SQLSmith 0.571 2.94 2.35
LearnedSQLGen 0.863 2.33 2.14
OmniSQL 0.716 1.97 5.58
SQL-Factory 0.637 4.10 7.94
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-driven random query generator that walks through SQL syntax trees; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of SQL queries.
For a fair comparison, we generate 10,000 SQL queries for the TPC-DS benchmark using each method. For OmniSQL, we adopt the original prompts provided by the authors and use both GPT-4o and Qwen2.5-Coder-14B for query generation. The proportion of queries generated by each model is kept consistent with the generation and expansion ratio used in SQL-Factory. We evaluate the diversity of generated queries across multiple evaluation metrics, including our hybrid similarity and vendi score.
Hybrid Similarity. We use our proposed hybrid similarity metric as described in Section 4.1.2 to measure the average pairwise similarity between queries. This metric integrates token-level, structural-level, and embedding-level similarity signals, providing a robust measure of redundancy in query content and structure. Lower scores indicate higher diversity. As shown in Table 2, SQL-Factory achieves a balanced similarity score of 0.637, significantly lower than LearnedSQLGen (0.863) and OmniSQL (0.688), suggesting that it avoids over-repetition while maintaining meaningful structure. Besides, the low similarity score for SQLSmith stems largely from its use of excessive syntactic randomness. It frequently inserts arbitrary user-defined functions, opaque predicates, and irrelevant clauses that contribute to lexical variance but not to meaningful semantic or structural diversity.
Vendi Score. The Vendi Score (Friedman and Dieng, 2023) is a recently proposed metric for evaluating diversity based on the exponential of the entropy of the eigenvalues of a similarity matrix computed from the data. It does not rely on predefined labels or reference distributions, making it especially suitable for open-ended generation tasks. We utilize Vendi Score to assess the diversity of our generated datasets with two embedding backbones: SimCSE (Gao et al., 2021), a general-purpose sentence encoder trained via contrastive learning, and SQL-Encoder (Pourreza et al., 2024), a specialized model pretrained for SQL embeddings. Higher Vendi scores indicate greater overall diversity in the embedding space. As shown in Table 2, SQL-Factory achieves the highest Vendi Score under both encoders. In contrast, SQLSmith demonstrates low similarity but also fails to produce coherent queries. LearnedSQLGen suffers from high redundancy due to its pattern imitation and optimization-driven training. Besides, OmniSQL is limited by its monolithic generation strategy.
5.2.4. Schema-Aware Query Allocation Analysis
((a))
((b))
Figure 7. Query and Complexity Distributions in OmniSQL.
((a))
((b))
Figure 8. Query and Complexity Distributions in SQL-Factory.
To evaluate whether SQL-Factory enables schema-aware query planning, we compare it with OmniSQL (Li et al., 2025a) across two key metrics: the number of queries for each table and the structural complexity of each table. We define a table’s complexity score as the sum of its number of columns and twice the number of foreign keys, which reflects both structural richness and relational connectivity. Figure 7 and Figure 8 illustrate these distributions for OmniSQL and SQL-Factory, respectively.
Figure 7(a) shows that exhibits a highly skewed distribution, where a small number of tables dominate the workload while most tables receive very few queries. However, as shown in Figure Figure 7(b), the distribution of table complexity in the same benchmark is much more balanced, with the majority of tables falling in a moderate complexity range. This misalignment suggests that OmniSQL does not account for schema structure when allocating queries, which may result in under-representation of complex tables.
In contrast, SQL-Factory produces a query distribution that more closely matches schema complexity. Figure 8(a) shows that most tables receive a moderate number of queries, and this pattern aligns well with the table complexity distribution in Figure 8(b), where moderately complex tables are also the most common. This consistency indicates that SQL-Factory can allocate queries in a way that reflects the underlying schema properties, supporting schema-aware planning and improving the quality of the workload.
5.2.5. Generation Stopping Criteria
((a))
((b))
((c))
((d))
((e))
Figure 9. Evolution of Expansion Ratio and Overall Similarity as the Number of Synthesized SQL Queries Increases.
Although SQL-Factory is capable of generating an unbounded number of SQL queries for any benchmark, it is also important to ensure high overall data quality. To this end, we design a generation termination policy that avoids excessive accumulation of redundant or low-diversity queries.
To determine when to terminate for each benchmark, we monitor two key metrics: (1) the expansion ratio, defined as the proportion of all queries in the SQL Pool that originate from the Expansion Team, and (2) the average hybrid similarity, which reflects query redundancy in the overall pool.
As illustrated in Figure 9, the system is initialized by the Generation Team, which produces a small number of diverse seed queries. After this warm-up phase, the Expansion Team begins to take over, resulting in a rise in both the expansion ratio and similarity across the SQL Pool. This indicates that a growing number of queries are being synthesized based on previously explored patterns.
To maintain overall query diversity, the Management Agent periodically switches back to the Generation Team. These transitions manifest as synchronized dips in both expansion ratio and similarity on the curve, reflecting the injection of more novel queries. This alternating pattern of generation and expansion continues iteratively, with diminishing gains over time. Once the Expansion Team becomes dominant and the similarity curve shows a sustained upward trend, it indicates that the SQL Pool has reached semantic saturation, prompting us to terminate the synthesis process.
For example, after synthesizing approximately 50,000 queries for TPC-DS, we observe that the Expansion Team begins to dominate the process. Specifically, the expansion ratio rapidly decrease while the average similarity among queries steadily rising. This trend indicates that the SQL Pool has become saturated. Continuing the process beyond this point leads to diminishing returns in diversity and semantic coverage. Therefore, we empirically adopt this saturation point as the stopping criterion for each benchmark.
5.3. Text-to-SQL Task
Table 3. Results on Text-to-SQL Task.
Method Spider Bird
Executability Accuracy Executability Accuracy
Avg Avg Simple Moderate Challenging Avg Simple Moderate Challenging Avg
Codes-1B 90.43 68.09 89.30 81.03 83.45 86.25 42.38 26.51 22.07 35.66
Codes-3B 94.97 75.05 92.32 85.34 88.28 89.83 53.30 33.19 28.97 44.92
Qwen2.5-Coder-1.5B 85.69 69.15 63.24 49.78 46.21 57.56 34.92 19.83 15.86 28.55
Qwen2.5-Coder-3B 92.84 76.21 73.51 63.15 56.55 68.77 43.78 29.74 17.93 37.09
Qwen2.5-Coder-1.5B+SQL-Factory 97.20 76.50 87.89 74.57 66.21 81.81 50.92 26.94 19.31 40.68
Qwen2.5-Coder-3B+SQL-Factory 97.29 79.69 92.97 82.11 75.86 88.07 55.35 33.62 26.90 46.09
To further validate the practical value of the SQL-Factory framework, we evaluate its effectiveness in improving performance on downstream Text-to-SQL tasks.
5.3.1. Dataset Construction
Since SQL-Factory is primarily designed to synthesize executable SQL queries for given database schema from scratch, direct application to Text-to-SQL tasks requires transforming these SQLs into natural language (NL) descriptions. This enables the construction of ¡NL, SQL¿ training pairs for supervised learning.
Instead of generating SQL from NL, which often suffers from semantic mismatch and low exact match accuracy, we adopt a reverse formulation: converting SQL to NL. Recent studies (Wu et al., 2021; Guo et al., 2018; Zhang et al., 2023) have shown that LLMs are more reliable in generating faithful natural language descriptions for given SQL. This makes the SQL-to-NL approach a robust and efficient means for constructing high-quality data. To this end, we develop a multi-stage pipeline for SQL-to-NL conversion with consistency checking and optional SQL refinement. The steps are as follows:
(1) Similar Example Retrieval. To improve translation quality and ensure semantic alignment, we first retrieve top-$k$ similar examples (with $k$=5 in the experiment) from training set of BIRD or Spider using Retrieval Tool proposed in Section 4.1.2. These examples are used to construct the in-context prompt to teach model how to write fluent, semantically faithful questions from SQL. This retrieval-augmented generation ensures structural and linguistic alignment between training and inference distributions.
(2) SQL-to-NL Conversion. We then adopt a prompt-based method to translate SQL queries into natural language using a powerful instruction-tuned language model (Qwen2.5-Coder-14B(Hui et al., 2024)). Given the database schema, SQL query and a few-shot template, the model generates a corresponding natural language question that accurately reflects the intent of the SQL. Additionally, to ensure consistency with the BIRD dataset format, each generated NL question is accompanied by an external knowledge explanation.
(3) Semantic Consistency Verification. To further ensure the correctness of the constructed ¡NL, SQL¿ pairs, we introduce an LLM-based verifier to perform semantic consistency checking. This component is critical for filtering out misaligned instances where the generated natural language question does not faithfully reflect the intent or logic of the corresponding SQL query. Specifically, the verifier takes as input the SQL query and its corresponding natural language question, and is prompted to return a binary decision indicating whether the pair is semantically aligned. To improve reasoning quality and reduce false positives, we adopt a Chain-of-Thought (CoT) prompting strategy. Only those ¡NL, SQL¿ pairs that pass this verification step are retained, ensuring high semantic fidelity in the final dataset.
(4) SQL Simplification. In case where the SQL query contains redundant columns or sub-clauses not reflected in the NL intent, we use a follow-up prompt to refine the SQL accordingly. This ensures alignment between the question intent and query logic. All simplified SQLs are passed again through the verifier to ensure final consistency.
This multi-stage pipeline enables us to automatically transform SQL-only corpora generated by SQL-Factory into a large-scale, high-quality Text-to-SQL training set, effectively bridging the gap between unsupervised SQL generation and downstream tasks.
5.3.2. Model Training
We fine-tune two backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
5.3.3. Result and Analysis
To ensure fairness, we adopt the official prompt templates for each model. For codes models, we use the released prompt from the original repository but remove schema linking and value retrieval modules to match our setup. For all Qwen-family models, we adopt the default instruction prompt used by Qwen2.5-Coder (Hui et al., 2024). Each model is evaluated on two benchmarks: Spider and BIRD and the evaluation metrics include both executability and execution accuracy.
The results are shown in Table 3. After fine-tuning with the SQL-Factory dataset, both Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B exhibit substantial performance improvements. On Spider, Qwen2.5-Coder-1.5B+SQL-Factory achieves an executability of 97.20% and accuracy of 76.50%, while Qwen2.5-Coder-3B+SQL-Factory reaches 97.29% executability and 79.69% accuracy. Both outperform their corresponding base models by a large margin, demonstrate that SQL-Factory effectively boosts model performance even without increasing model size. On BIRD, which contains more complex and compositional queries, fine-tuned models again show superior performance. Qwen2.5-Coder-3B+SQL-Factory achieves 88.07% in executability and 46.09% in accuracy, clearly outperforming all baseline models of similar scale. These improvements are especially pronounced in the “Moderate” and “Challenging” categories, confirming the effectiveness of SQL-Factory in providing structurally rich and semantically aligned training data.
Overall, these findings show that with access to SQL-Factory, small and mid-sized models can achieve strong performance in Text-to-SQL tasks. This highlights SQL-Factory’s ability to deliver diverse and executable training examples that bridge the performance gap without requiring massive model sizes.
5.4. Query Clustering Task
Table 4. Performance on Query Clustering Task.
Methods IIT Bombay UB Exam PocketData
Aouiche (Aouiche et al., 2006) 0.577 0.923 0.893
Aligon (Aligon et al., 2014) 0.535 0.799 0.898
Makiyama (Makiyama et al., 2015) 0.665 0.897 0.879
BERT-Base (Devlin et al., 2019) 0.524 0.769 0.773
BERT-Large (Devlin et al., 2019) 0.637 0.846 0.777
PreQR (Tang et al., 2022) 0.387 0.622 0.710
BERT-Base+SQL-Factory 0.376 0.558 0.656
BERT-Large+SQL-Factory 0.350 0.549 0.638
To further evaluate the utility of the SQL queries generated by our SQL-Factory, we investigate their effectiveness in improving SQL representation for query clustering tasks.
We fine-tune both BERT-Base and BERT-Large models using all SQL queries synthesized by SQL-Factory, aiming to enhance their capability in capturing SQL-level semantics and improve performance in downstream clustering tasks. The training process employs contrastive learning, a method that leverages positive and negative sample pairs to enhance model understanding. Positive pairs are selected via the Retrieval Tool introduced in Section 4.1.2. Specifically, for each anchor query, we retrieve its top-20 most similar queries based on our hybrid similarity metric and treat them as positive samples. All remaining queries within the same batch serve as negative samples. This in-batch negative sampling strategy ensures scalability and efficiency, as it avoids the need to explicitly mine hard negatives from the full dataset while still providing sufficient contrastive signal.
After training, we apply the fine-tuned models to encode all queries from the clustering benchmarks. Pairwise distances between queries are computed using cosine similarity in the embedding space. We evaluate clustering quality using the BetaCV (Mohammed and Wagner, 2014) metric, which measures intra-cluster compactness relative to inter-cluster distance. Specifically, a lower BetaCV score indicates better clustering performance, as it reflects that queries within the same cluster are more tightly grouped, while those from different clusters are more widely separated.
Table 4 presents the BetaCV scores across all models and benchmarks. Traditional rule-based methods such as Aouiche, Aligon, and Makiyama rely on handcrafted SQL features or syntax trees, and often fail to capture deep semantic similarities across structurally different but semantically related queries. PreQR improves upon this by leveraging a pretrained language model, but it lacks task-specific supervision and may underfit domain-specific patterns. Meanwhile, untuned BERT models benefit from general semantic priors but are still agnostic to SQL-specific structure and vocabulary.
By contrast, our contrastively fine-tuned BERT models significantly outperform all baselines. Both versions of the BERT model have achieved the best performance on these benchmarks after training. These results validate the effectiveness of SQL-Factory in producing high-quality, semantically diverse SQL queries that can be used as training signals for representation learning.
6. Conclusion
In this paper, we propose SQL-Factory, a multi-agent framework designed to generate high-quality and large-scale SQL queries. By organizing the system into three collaborative teams for generation, expansion, and management, SQL-Factory achieves a balance between diversity, efficiency, and cost. The framework leverages powerful language models to explore novel query structures, while lightweight local models efficiently scale promising patterns. Through real-time evaluation and schema-aware planning, SQL-Factory dynamically adjusts generation focus based on coverage gaps and schema complexity, ensuring that structurally rich tables receive sufficient attention. We demonstrate the effectiveness of SQL-Factory by applying it to multiple benchmarks, generating over 300,000 high-quality queries with minimal cost. Extensive experiments show that the generated queries significantly enhance downstream tasks such as Text-to-SQL and query clustering, confirming SQL-Factory as a practical solution for building large-scale, high-utility SQL corpora.